# A Diffusion-Based Approach for Simulating Forward-in-Time State-Dependent Speciation and Extinction Dynamics

### v1

@@ 3.2 Multiple rate scenarios lead to the same stationary state frequencies
$\hat{\Pi}_{A}=\frac{1}{3},\hat{\Pi}_{B}=\frac{1}{3},\hat{\Pi}_{AB}=\frac{1}{3}.$
@@
Figure 7: The expected range state dynamics over [0,250] time interval and over 100 trajectories for the two-region GeoSSE model as described in Example. 5. Each process is simulated under the following initial state frequencies and rate parameters according to Eq. (99): (Left panel) $\Pi_{A}(0)=\Pi_{B}(0)=0.45,\Pi_{AB}(0)=0.1,w_{A}=0.090,e_{A}=0.176,w_{B}\approx 0,e_{B}=0.008,d_{AB}=0.015,d_{BA}=0.173,b^{A}_{B}=0.004$ ; (Right panel) $\Pi_{A}(0)=0.1,\Pi_{B}(0)=\Pi_{AB}(0)=0.45,w_{A}=0.160,e_{A}=0.315,w_{B}=0.002,e_{B}=0.009,d_{AB}=0.014,d_{BA}=0.310,b^{A}_{B}=0.001$. In both panels, $\mathbb{E}(\hat{\Pi}_{A})\rightarrow\frac{1}{3},\mathbb{E}(\hat{\Pi}_{B})\rightarrow\frac{1}{3},\mathbb{E}(\hat{\Pi}_{AB})\rightarrow\frac{1}{3}$. Using Lemma 5, we confirm that these expected stationary frequencies from simulations converge to the theoretical, and true stationary frequencies given these sets of rates. Furthermore, using the procedure described in Section 2.8 with $\epsilon=10^{-9}$, we found that the stationary frequencies are reached at: $t^{*}_{A}=114.114,\>t^{*}_{B}=111.862,t^{*}_{AB}=102.603$ (Left panel); $t^{*}_{A}=76.827,\>t^{*}_{B}=75.576,t^{*}_{AB}=70.320$ (Right panel).
@@
Figure 8: The expected range state dynamics over [0,60] time interval and over 100 trajectories for the two-region GeoSSE model as described in Example. 5. Each process is simulated under the following initial state frequencies and rate parameters according to Eq. (100): (Left panel) $\Pi_{A}(0)=\Pi_{B}(0)=0.45,\Pi_{AB}(0)=0.1,w_{A}=0.107,e_{A}=0.309,w_{B}=0.008,e_{B}=0.008,d_{AB}=0.001,d_{BA}=0.405,b^{A}_{B}=0.089$ ; (Right panel) $\Pi_{A}(0)=0.1,\Pi_{B}(0)=\Pi_{AB}(0)=0.45,w_{A}=0.049,e_{A}=0.470,w_{B}=0.005,e_{B}=0.008,d_{AB}=0.006,d_{BA}=0.843,b^{A}_{B}=0.371$. In both panels, $\mathbb{E}(\hat{\Pi}_{A})\rightarrow\frac{1}{3},\mathbb{E}(\hat{\Pi}_{B})\rightarrow\frac{1}{3},\mathbb{E}(\hat{\Pi}_{AB})\rightarrow\frac{1}{3}$. Using Lemma 5, we confirm that these expected stationary frequencies from simulations converge to the theoretical, and true stationary frequencies given these sets of rates. Furthermore, using the procedure described in Section 2.8 with $\epsilon=10^{-9}$, we found that the stationary frequencies are reached at: $t^{*}_{A}=53.153,\>t^{*}_{B}=51.952,t^{*}_{AB}=48.048$ (Left panel); $t^{*}_{A}=30.781,\>t^{*}_{B}=30.330,t^{*}_{AB}=28.378$ (Right panel).
@@ 3.3 Comparing our method of computing stationary state frequencies with existing literature
our method gives $\hat{\Pi}_{A}\approx 0.057,\>\hat{\Pi}_{B}\approx 0.506,\>\hat{\Pi}_{AB}\approx 0.437$ while the projection matrix approach implemented in diversitree returns $\hat{\Pi}_{A}\approx 0.055,\>\hat{\Pi}_{B}\approx 0.490,\>\hat{\Pi}_{AB}\approx 0.455$. Another example using the following rate parameters,
@@
we have $\hat{\Pi}_{A}\approx 0.0996,\>\hat{\Pi}_{B}\approx 0.0996,\>\hat{\Pi}_{AB}\approx 0.8008$ while the other method produces $\hat{\Pi}_{A}\approx 0.0997,\>\hat{\Pi}_{B}\approx 0.0997,\>\hat{\Pi}_{AB}\approx 0.8006$.
@@ 4 Discussion and Conclusion
In our paper, we have constructed a general framework using diffusion processes to study state dynamics over time from a general state-dependent speciation and extinction model with both anagenetic and cladogenetic state transitions, making it suitable for studying members of the ClaSSE model family [9, 26, 10, 8]. We have applied this framework under various diversification scenarios for the GeoSSE model [10], a special case of the ClaSSE model, as described in Sections 2.4-2.5. Our framework can easily be applied to other discrete state-dependent diversification models, such as simpler BiSSE and MuSSE models [25, 7] and Markovian Binary Tree (MBT) models [18, 12, 31]. Through simulations and statistical analyses, we have shown that state dynamics simulated under diffusion-based approach and tree-based approach are comparable (Figs. 3-6, Table 1). We also show, using a statistical test, that our diffusion framework offers a good approximation for the mean of state counts. This result allows one to understand how data generating process i.e. rate parameters from a diversification model can explain observed state patterns without using phylogenetic information. For inferring rates using empirical state data at present, this diffusion-based approach to simulate state dynamics could be treated as a way to validate whether rates estimated from biological datasets using phylogenetic methods are sensible.
Moreover, in Sections 2.6-2.7, we have derived theoretical results to deduce the expected state frequencies generated by a set of rates, and what possible rates will generate a given set of expected state frequencies. These results are generalizable to accommodate a system having more states, and provide an alternative way to validate the correctness of SSE simulation and inference methods. Additionally, in Section 2.8, we described a procedure to compute the minimum time for an SSE process to reach stationarity in its state frequencies. We have applied these results for a 2-region GeoSSE model. As seen in Figs. 7-8, we showed that there exist multiple different rate scenarios that can lead to the same stationary behaviour of state pattern. Our framework also creates an alternative mathematical approach to tree-based models that could help establish conditions for which SSE model parameters are and are not identifiable.
@@ 8.1 Proof of Lemma 1
From Eq. (1), we compute $g^{\prime}(N)$ and $g"(N)$,
@@ 8.2 Proof of Lemma 2
By definition,
$\mu_{i}$ $=$ $\lim_{\Delta t\rightarrow 0}\frac{1}{\Delta t}\{\mathbb{E}(N_{i}(t+\Delta t))-N_{i}\Delta t\},$
$\sigma_{i}^{2}$ $=$ $\lim_{\Delta t\rightarrow 0}\frac{1}{\Delta t}\{\mathbb{E}\left(N_{i}^{2}(t+\Delta t)\right)-2N_{i}\mathbb{E}\left(N_{i}(t+\Delta t)\right)+N_{i}^{2}\Delta t\}.$
@@
$=$ $N_{i}\Delta t+\left(\mathbb{P}_{i}^{+}-\mathbb{P}_{i}^{-}\right)\Delta t,$
@@ 8.3 Proof of Lemma 3
$=$ $\frac{1-\Pi_{i}}{N}\mu_{i}+\sum_{\begin{subarray}{c}j\in S\\ j\neq i\end{subarray}}\frac{-\Pi_{i}}{N}\mu_{j}-\frac{1-\Pi_{i}}{N^{2}}\sigma_{i}^{2}+\sum_{\begin{subarray}{c}j\in S\\ j\neq i\end{subarray}}\frac{\Pi_{i}}{N^{2}}\sigma_{j}^{2}$ (119)
$=$ $\left(\frac{1-\Pi_{i}}{N}\right)\left(\mu_{i}-\frac{\sigma_{i}^{2}}{N}\right)+\sum_{\begin{subarray}{c}j\in S\\ j\neq i\end{subarray}}\frac{\Pi_{i}}{N}\left(-\mu_{j}+\frac{\sigma_{j}^{2}}{N}\right)$ (122)
$=$ $\frac{1}{N}\left(\mu_{i}-\frac{\sigma^{2}_{i}}{N}\right)+\frac{\Pi_{i}}{N}\sum_{j\in S}\left(-\mu_{j}+\frac{\sigma_{j}^{2}}{N}\right),$ (123)
@@
$=$ $\left(\frac{1-\Pi_{i}}{N}\right)^{2}\sigma_{i}^{2}+\sum_{\begin{subarray}{c}j\in S\\ j\neq i\end{subarray}}\left(\frac{\Pi_{i}}{N}\right)^{2}\sigma_{j}^{2}$ (129)
$=$ $\left(\frac{\sigma_{i}}{N}\right)^{2}(1-2\Pi_{i})+\left(\frac{\Pi_{i}}{N}\right)^{2}\sum_{j\in S}\sigma_{j}^{2},$ (130)
@@ 8.4 Proof of Lemma 5
$=$ $\left[w_{A}\left(\Pi_{A}+\Pi_{AB}\right)+\Pi_{AB}b^{A}_{B}+e_{B}\Pi_{AB}\right]-\left[d_{AB}\Pi_{A}+e_{A}\Pi_{A}\right]$
$=$ $\Pi_{A}\left(w_{A}-e_{A}-d_{AB}\right)+\Pi_{AB}\left(w_{A}+b^{A}_{B}+e_{B}\right).$
@@
$=$ $\left[w_{B}\left(\Pi_{B}+\Pi_{AB}\right)+\Pi_{AB}b^{A}_{B}+e_{A}\Pi_{AB}\right]-\left[d_{BA}\Pi_{B}+e_{B}\Pi_{B}\right]$
$=$ $\Pi_{B}\left(w_{B}-e_{B}-d_{BA}\right)+\Pi_{AB}\left(w_{B}+b^{A}_{B}+e_{A}\right).$
@@
$=$ $\left[d_{AB}\Pi_{A}+d_{BA}\Pi_{B}\right]-\left[b^{A}_{B}\Pi_{AB}+(e_{A}+e_{B})\Pi_{AB}\right]$
$=$ $\left(d_{AB}\Pi_{A}+d_{BA}\Pi_{B}\right)-\Pi_{AB}\left(b^{A}_{B}+e_{A}+e_{B}\right).$
@@
$\frac{d\Pi_{A}}{dt}$ $=$ $\Pi_{A}\left(w_{A}-e_{A}-d_{AB}\right)+\Pi_{AB}\left(w_{A}+b^{A}_{B}+e_{B}\right),$ (134)
$\frac{d\Pi_{B}}{dt}$ $=$ $\Pi_{B}\left(w_{B}-e_{B}-d_{BA}\right)+\Pi_{AB}\left(w_{B}+b^{A}_{B}+e_{A}\right),$ (135)
$\frac{d\Pi_{AB}}{dt}$ $=$ $\left(d_{AB}\Pi_{A}+d_{BA}\Pi_{B}\right)-\Pi_{AB}\left(b^{A}_{B}+e_{A}+e_{B}\right).$ (136)
given initial state frequencies $\Pi_{A}(0)=\Pi_{A}^{0},\Pi_{B}(0)=\Pi_{B}^{0},\Pi_{AB}(0)=\Pi_{AB}^{0}$.
However, since $\hat{\Pi}_{A}+\hat{\Pi}_{B}+\hat{\Pi}_{AB}=1$, we can always derive $\hat{\Pi}_{AB}$ using $\hat{\Pi}_{A}$ and $\hat{\Pi}_{B}$. Therefore, we want to solve the following system instead.
$\frac{d\Pi_{A}}{dt}$ $=$ $\Pi_{A}\left(w_{A}-e_{A}-d_{AB}\right)+\Pi_{AB}\left(w_{A}+b^{A}_{B}+e_{B}\right),$ (137)
$\frac{d\Pi_{B}}{dt}$ $=$ $\Pi_{B}\left(w_{B}-e_{B}-d_{BA}\right)+\Pi_{AB}\left(w_{B}+b^{A}_{B}+e_{A}\right).$ (138)
Since $\Pi_{A}(t)+\Pi_{B}(t)+\Pi_{AB}(t)=1$, we have,
$\frac{d\Pi_{A}}{dt}$ $=$ $\Pi_{A}\left(-e_{A}-d_{AB}-b^{A}_{B}-e_{B}\right)-\Pi_{B}\left(w_{A}+b^{A}_{B}+e_{B}\right)+\left(w_{A}+b^{A}_{B}+e_{B}\right),$ (139)
$\frac{d\Pi_{B}}{dt}$ $=$ $\Pi_{B}\left(-e_{B}-d_{BA}-b^{A}_{B}-e_{A}\right)-\Pi_{A}\left(w_{B}+b^{A}_{B}+e_{A}\right)+\left(w_{B}+b^{A}_{B}+e_{A}\right).$ (140)
@@
$\bm{\Pi^{{}^{\prime}}}=\begin{bmatrix}\Pi_{A}^{{}^{\prime}}\\ \Pi_{B}^{{}^{\prime}}\end{bmatrix},\>\bm{\Pi}=\begin{bmatrix}\Pi_{A}\\ \Pi_{B}\end{bmatrix},\>M=\begin{bmatrix}-\left(e_{A}+d_{AB}+b^{A}_{B}+e_{B}\right)&-\left(w_{A}+b^{A}_{B}+e_{B}\right)\\ -\left(w_{B}+b^{A}_{B}+e_{A}\right)&-\left(e_{B}+d_{BA}+b^{A}_{B}+e_{A}\right)\end{bmatrix},\>\bm{r}=\begin{bmatrix}w_{A}+b^{A}_{B}+e_{B}\\ w_{B}+b^{A}_{B}+e_{A}\end{bmatrix}.$ (150)
@@
$R=\sqrt{4\left(b^{A}_{B}\right)^{2}+4\left(b^{A}_{B}e_{A}+b^{A}_{B}e_{B}+b^{A}_{B}w_{A}+b^{A}_{B}w_{B}\right)+4\left(e_{A}e_{B}+e_{A}w_{A}+e_{B}w_{B}+w_{A}w_{B}\right)-2d_{AB}d_{BA}+\left(d_{AB}^{2}+d_{BA}^{2}\right)}.$
@@
$\begin{bmatrix}\hat{\Pi}_{A}\\ \hat{\Pi}_{B}\end{bmatrix}$ $=$ $\begin{bmatrix}K_{1}\\ K_{2}\end{bmatrix}.$ (183)
$\hat{\Pi}_{A}$ and $\hat{\Pi}_{B}$ from Eq. (183) are the stationary frequencies for state $A$ and $B$, respectively, as shown in Lemma 5.
@@
$C_{1}\begin{bmatrix}-\frac{1}{2\left(b^{A}_{B}+e_{A}+w_{B}\right)}\left(-d_{AB}+d_{BA}-R\right)\\[10.0pt] 1\end{bmatrix}+C_{2}\begin{bmatrix}-\frac{1}{2\left(b^{A}_{B}+e_{A}+w_{B}\right)}\left(-d_{AB}+d_{BA}+R\right)\\[10.0pt] 1\end{bmatrix}+\begin{bmatrix}K_{1}\\[10.0pt] K_{2}\end{bmatrix}=\begin{bmatrix}\Pi_{A}^{0}\\[10.0pt] \Pi_{B}^{0}\end{bmatrix}.$ (192)

### v2

@@ 3.2 Multiple rate scenarios lead to the same stationary state frequencies
$\hat{\Pi}_{\{A\}}=\frac{1}{3},\hat{\Pi}_{\{B\}}=\frac{1}{3},\hat{\Pi}_{\{A,B\}}=\frac{1}{3}.$
@@
Figure 9: The expected range state dynamics over [0,250] time interval and over 100 trajectories for the two-region GeoSSE model as described in Example. 5. Each process is simulated under the following initial state frequencies and rate parameters according to Eq. (34): (Left panel) $\Pi_{\{A\}}(0)=\Pi_{\{B\}}(0)=0.45,\Pi_{\{A,B\}}(0)=0.1,w_{A}=0.090,e_{A}=0.176,w_{B}\approx 0,e_{B}=0.008,d_{AB}=0.015,d_{BA}=0.173,b^{A}_{B}=0.004$ ; (Right panel) $\Pi_{\{A\}}(0)=0.1,\Pi_{\{B\}}(0)=\Pi_{\{A,B\}}(0)=0.45,w_{A}=0.160,e_{A}=0.315,w_{B}=0.002,e_{B}=0.009,d_{AB}=0.014,d_{BA}=0.310,b^{A}_{B}=0.001$. In both panels, $\mathbb{E}(\hat{\Pi}_{\{A\}})\rightarrow\frac{1}{3},\mathbb{E}(\hat{\Pi}_{\{B\}})\rightarrow\frac{1}{3},\mathbb{E}(\hat{\Pi}_{\{A,B\}})\rightarrow\frac{1}{3}$. Using Lemma 5, we confirm that these expected stationary frequencies from simulations converge to the theoretical, and true stationary frequencies given these sets of rates. Furthermore, using the procedure described in Section 2.8 with $\epsilon=10^{-9}$, we found that the stationary frequencies are reached at: $t^{*}_{A}=114.114,\>t^{*}_{B}=111.862,t^{*}_{AB}=102.603$ (Left panel); $t^{*}_{A}=76.827,\>t^{*}_{B}=75.576,t^{*}_{AB}=70.320$ (Right panel)
@@
Figure 10: The expected range state dynamics over [0,60] time interval and over 100 trajectories for the two-region GeoSSE model as described in Example. 5. Each process is simulated under the following initial state frequencies and rate parameters according to Eq. (35): (Left panel) $\Pi_{\{A\}}(0)=\Pi_{\{B\}}(0)=0.45,\Pi_{\{A,B\}}(0)=0.1,w_{A}=0.107,e_{A}=0.309,w_{B}=0.008,e_{B}=0.008,d_{AB}=0.001,d_{BA}=0.405,b^{A}_{B}=0.089$ ; (Right panel) $\Pi_{\{A\}}(0)=0.1,\Pi_{\{B\}}(0)=\Pi_{\{A,B\}}(0)=0.45,w_{A}=0.049,e_{A}=0.470,w_{B}=0.005,e_{B}=0.008,d_{AB}=0.006,d_{BA}=0.843,b^{A}_{B}=0.371$. In both panels, $\mathbb{E}(\hat{\Pi}_{\{A\}})\rightarrow\frac{1}{3},\mathbb{E}(\hat{\Pi}_{\{B\}})\rightarrow\frac{1}{3},\mathbb{E}(\hat{\Pi}_{\{A,B\}})\rightarrow\frac{1}{3}$. Using Lemma 5, we confirm that these expected stationary frequencies from simulations converge to the theoretical, and true stationary frequencies given these sets of rates. Furthermore, using the procedure described in Section 2.8 with $\epsilon=10^{-9}$, we found that the stationary frequencies are reached at: $t^{*}_{A}=53.153,\>t^{*}_{B}=51.952,t^{*}_{AB}=48.048$ (Left panel); $t^{*}_{A}=30.781,\>t^{*}_{B}=30.330,t^{*}_{AB}=28.378$ (Right panel)
@@ 3.3 Comparing our method of computing stationary state frequencies with existing literature
our method gives $\hat{\Pi}_{\{A\}}\approx 0.057,\>\hat{\Pi}_{\{B\}}\approx 0.506,\>\hat{\Pi}_{\{A,B\}}\approx 0.437$ while the projection matrix approach implemented in diversitree returns $\hat{\Pi}_{\{A\}}\approx 0.055,\>\hat{\Pi}_{\{B\}}\approx 0.490,\>\hat{\Pi}_{\{A,B\}}\approx 0.455$. Another example using the following rate parameters,
@@
we have $\hat{\Pi}_{\{A\}}\approx 0.0996,\>\hat{\Pi}_{\{B\}}\approx 0.0996,\>\hat{\Pi}_{\{A,B\}}\approx 0.8008$ while the other method produces $\hat{\Pi}_{\{A\}}\approx 0.0997,\>\hat{\Pi}_{\{B\}}\approx 0.0997,\>\hat{\Pi}_{\{A,B\}}\approx 0.8006$.
@@ 4 Discussion and Conclusion
In our paper, we have constructed a general framework using diffusion processes to study state dynamics over time from a general state-dependent speciation and extinction model with both anagenetic and cladogenetic state transitions, making it suitable for studying members of the ClaSSE model family [10, 26, 11, 9]. We have applied this framework under various diversification scenarios for the GeoSSE model [11], a special case of the ClaSSE model, as described in Sections 2.4-2.5. Our framework can easily be applied to other discrete state-dependent diversification models, such as simpler BiSSE and MuSSE models [25, 8] and Markovian Binary Tree (MBT) models [18, 13, 31]. Through simulations and statistical analyses, we have shown that state dynamics simulated under diffusion-based approach and tree-based approach are comparable, given that we start the simulations with relative large clade size (Figs. 5-8, Table 1). We also obtain good agreement between diffusion-based and tree-based simulations when beginning the process with a single species in random state, after applying a model-based correction procedure (Appendix 7.6, Fig. 12). We also show, using a statistical test, that our diffusion framework offers a good approximation for the mean of state counts. This result allows one to understand how data generating process i.e. rate parameters from a diversification model can explain observed state patterns without using phylogenetic information. For inferring rates using empirical state data at present, this diffusion-based approach to simulate state dynamics could be treated as a way to validate whether rates estimated from biological datasets using phylogenetic methods are sensible.
Moreover, in Sections 2.6-2.7, we have derived theoretical results to deduce the expected state frequencies generated by a set of rates, and what possible rates will generate a given set of expected state frequencies. These results are generalizable to accommodate a system having more states, and provide an alternative way to validate the correctness of SSE simulation and inference methods. Additionally, in Section 2.8, we described a procedure to compute the minimum time for an SSE process to reach stationarity in its state frequencies. We have applied these results for a 2-region GeoSSE model. As seen in Figs. 9-10, we showed that there exist multiple different rate scenarios that can lead to the same stationary behaviour of state pattern. Our framework also creates an alternative mathematical approach to tree-based models that could help establish conditions for which SSE model parameters are and are not identifiable.
@@ 6 Acknowledgements
Data availability
Conflict of interest
The authors declare that they have no conflict of interest.
@@ 7.1 Proof of Lemma 1
From Eq. (1), we compute $g^{\prime}(N)$ and $g^{\prime\prime}(N)$,
@@ 7.2 Proof of Lemma 2
By definition, following Eqs. 1.2–1.3 of Chapter 5 in [17], we define the infinitesimal mean $\mu_{i}$ and infinitesimal variance $\sigma_{i}^{2}$ of the stochastic process $\{N_{i}(t):t>0\}$ as follows,
$\mu_{i}$ $=$ $\lim_{\Delta t\rightarrow 0}\frac{1}{\Delta t}\mathbb{E}\left(N_{i}(t+\Delta t)-N_{i}(t)|N_{i}(t)=N_{i}\right)$
$=$ $\lim_{\Delta t\rightarrow 0}\frac{1}{\Delta t}\{\mathbb{E}(N_{i}(t+\Delta t))-\mathbb{E}(N_{i}(t)|N_{i}(t)=N_{i})\}$
$=$ $\lim_{\Delta t\rightarrow 0}\frac{1}{\Delta t}\{\mathbb{E}(N_{i}(t+\Delta t))-N_{i}\},$
$\sigma_{i}^{2}$ $=$ $\lim_{\Delta t\rightarrow 0}\frac{1}{\Delta t}\mathbb{E}\left((N_{i}(t+\Delta t)-N_{i}(t))^{2}|N_{i}(t)=N_{i}\right)$
$=$ $\lim_{\Delta t\rightarrow 0}\frac{1}{\Delta t}\mathbb{E}\left(N_{i}^{2}(t+\Delta t)-2N_{i}(t)N_{i}(t+\Delta t)+N_{i}^{2}(t)|N_{i}(t)\right)$
$=$ $\lim_{\Delta t\rightarrow 0}\frac{1}{\Delta t}\{\mathbb{E}\left(N_{i}^{2}(t+\Delta t)\right)-2N_{i}\mathbb{E}\left(N_{i}(t+\Delta t)\right)+N_{i}^{2}\}.$
@@
$=$ $N_{i}+\left(\mathbb{P}_{i}^{+}-\mathbb{P}_{i}^{-}\right)\Delta t,$
@@ 7.3 Proof of Lemma 3
$=$ $\frac{1-\Pi_{i}}{N}\mu_{i}+\sum_{\begin{subarray}{c}j\in S\\ j\neq i\end{subarray}}\frac{-\Pi_{i}}{N}\mu_{j}-\frac{1-\Pi_{i}}{N^{2}}\sigma_{i}^{2}+\sum_{\begin{subarray}{c}j\in S\\ j\neq i\end{subarray}}\frac{\Pi_{i}}{N^{2}}\sigma_{j}^{2}$
$=$ $\left(\frac{1-\Pi_{i}}{N}\right)\left(\mu_{i}-\frac{\sigma_{i}^{2}}{N}\right)+\sum_{\begin{subarray}{c}j\in S\\ j\neq i\end{subarray}}\frac{\Pi_{i}}{N}\left(-\mu_{j}+\frac{\sigma_{j}^{2}}{N}\right)$
$=$ $\frac{1}{N}\left(\mu_{i}-\frac{\sigma^{2}_{i}}{N}\right)+\frac{\Pi_{i}}{N}\sum_{j\in S}\left(-\mu_{j}+\frac{\sigma_{j}^{2}}{N}\right),$
@@
$=$ $\left(\frac{1-\Pi_{i}}{N}\right)^{2}\sigma_{i}^{2}+\sum_{\begin{subarray}{c}j\in S\\ j\neq i\end{subarray}}\left(\frac{\Pi_{i}}{N}\right)^{2}\sigma_{j}^{2}$
$=$ $\left(\frac{\sigma_{i}}{N}\right)^{2}(1-2\Pi_{i})+\left(\frac{\Pi_{i}}{N}\right)^{2}\sum_{j\in S}\sigma_{j}^{2},$
@@ 7.4 Proof of Lemma 5
$=$ $\left[w_{A}\left(\Pi_{\{A\}}+\Pi_{\{A,B\}}\right)+\Pi_{\{A,B\}}b^{A}_{B}+e_{B}\Pi_{\{A,B\}}\right]-\left[d_{AB}\Pi_{\{A\}}+e_{A}\Pi_{\{A\}}\right]$
$=$ $\Pi_{\{A\}}\left(w_{A}-e_{A}-d_{AB}\right)+\Pi_{\{A,B\}}\left(w_{A}+b^{A}_{B}+e_{B}\right).$
@@
$=$ $\left[w_{B}\left(\Pi_{\{B\}}+\Pi_{\{A,B\}}\right)+\Pi_{\{A,B\}}b^{A}_{B}+e_{A}\Pi_{\{A,B\}}\right]-\left[d_{BA}\Pi_{\{B\}}+e_{B}\Pi_{\{B\}}\right]$
$=$ $\Pi_{\{B\}}\left(w_{B}-e_{B}-d_{BA}\right)+\Pi_{\{A,B\}}\left(w_{B}+b^{A}_{B}+e_{A}\right).$
@@
$=$ $\left[d_{AB}\Pi_{\{A\}}+d_{BA}\Pi_{\{B\}}\right]-\left[b^{A}_{B}\Pi_{\{A,B\}}+(e_{A}+e_{B})\Pi_{\{A,B\}}\right]$
$=$ $\left(d_{AB}\Pi_{\{A\}}+d_{BA}\Pi_{\{B\}}\right)-\Pi_{\{A,B\}}\left(b^{A}_{B}+e_{A}+e_{B}\right).$
@@
$\frac{d\Pi_{\{A\}}}{dt}$ $=$ $\Pi_{\{A\}}\left(w_{A}-e_{A}-d_{AB}\right)+\Pi_{\{A,B\}}\left(w_{A}+b^{A}_{B}+e_{B}\right),$ (47)
$\frac{d\Pi_{\{B\}}}{dt}$ $=$ $\Pi_{\{B\}}\left(w_{B}-e_{B}-d_{BA}\right)+\Pi_{\{A,B\}}\left(w_{B}+b^{A}_{B}+e_{A}\right),$ (48)
$\frac{d\Pi_{\{A,B\}}}{dt}$ $=$ $\left(d_{AB}\Pi_{\{A\}}+d_{BA}\Pi_{\{B\}}\right)-\Pi_{\{A,B\}}\left(b^{A}_{B}+e_{A}+e_{B}\right).$ (49)
given initial state frequencies $\Pi_{\{A\}}(0)=\Pi_{A}^{0},\Pi_{\{B\}}(0)=\Pi_{B}^{0},\Pi_{\{A,B\}}(0)=\Pi_{AB}^{0}$.
However, since $\hat{\Pi}_{\{A\}}+\hat{\Pi}_{\{B\}}+\hat{\Pi}_{\{A,B\}}=1$, we can always derive $\hat{\Pi}_{\{A,B\}}$ using $\hat{\Pi}_{\{A\}}$ and $\hat{\Pi}_{\{B\}}$. Therefore, we want to solve the following system instead.
$\frac{d\Pi_{\{A\}}}{dt}$ $=$ $\Pi_{\{A\}}\left(w_{A}-e_{A}-d_{AB}\right)+\Pi_{\{A,B\}}\left(w_{A}+b^{A}_{B}+e_{B}\right),$ (50)
$\frac{d\Pi_{\{B\}}}{dt}$ $=$ $\Pi_{\{B\}}\left(w_{B}-e_{B}-d_{BA}\right)+\Pi_{\{A,B\}}\left(w_{B}+b^{A}_{B}+e_{A}\right).$ (51)
Since $\Pi_{\{A\}}(t)+\Pi_{\{B\}}(t)+\Pi_{\{A,B\}}(t)=1$, we have,
$\frac{d\Pi_{\{A\}}}{dt}$ $=$ $\Pi_{\{A\}}\left(-e_{A}-d_{AB}-b^{A}_{B}-e_{B}\right)-\Pi_{\{B\}}\left(w_{A}+b^{A}_{B}+e_{B}\right)+\left(w_{A}+b^{A}_{B}+e_{B}\right),$
$\frac{d\Pi_{\{B\}}}{dt}$ $=$ $\Pi_{\{B\}}\left(-e_{B}-d_{BA}-b^{A}_{B}-e_{A}\right)-\Pi_{\{A\}}\left(w_{B}+b^{A}_{B}+e_{A}\right)+\left(w_{B}+b^{A}_{B}+e_{A}\right).$
@@
$\bm{\Pi^{{}^{\prime}}}=\begin{bmatrix}\Pi_{\{A\}}^{{}^{\prime}}\\ \Pi_{\{B\}}^{{}^{\prime}}\end{bmatrix},\>\bm{\Pi}=\begin{bmatrix}\Pi_{\{A\}}\\ \Pi_{\{B\}}\end{bmatrix},\>M=\begin{bmatrix}-\left(e_{A}+d_{AB}+b^{A}_{B}+e_{B}\right)&-\left(w_{A}+b^{A}_{B}+e_{B}\right)\\ -\left(w_{B}+b^{A}_{B}+e_{A}\right)&-\left(e_{B}+d_{BA}+b^{A}_{B}+e_{A}\right)\end{bmatrix},\>$
$\bm{r}=\begin{bmatrix}w_{A}+b^{A}_{B}+e_{B}\\ w_{B}+b^{A}_{B}+e_{A}\end{bmatrix}.$ (55)
@@
$R=\sqrt{R_{1}+R_{2}},$
$R_{2}=-2d_{AB}d_{BA}+\left(d_{AB}^{2}+d_{BA}^{2}\right).$
@@
$\begin{bmatrix}\hat{\Pi}_{\{A\}}\\ \hat{\Pi}_{\{B\}}\end{bmatrix}$ $=$ $\begin{bmatrix}K_{1}\\ K_{2}\end{bmatrix}.$ (70)
$\hat{\Pi}_{\{A\}}$ and $\hat{\Pi}_{\{B\}}$ from Eq. (70) are the stationary frequencies for state $\{A\}$ and $\{B\}$, respectively, as shown in Lemma 5.
@@
$C_{1}\begin{bmatrix}-\frac{1}{2\left(b^{A}_{B}+e_{A}+w_{B}\right)}\left(-d_{AB}+d_{BA}-R\right)\\[10.0pt] 1\end{bmatrix}+C_{2}\begin{bmatrix}-\frac{1}{2\left(b^{A}_{B}+e_{A}+w_{B}\right)}\left(-d_{AB}+d_{BA}+R\right)\\[10.0pt] 1\end{bmatrix}+\begin{bmatrix}K_{1}\\[10.0pt] K_{2}\end{bmatrix}=\begin{bmatrix}\Pi_{A}^{0}\\[10.0pt] \Pi_{B}^{0}\end{bmatrix}.$